# Comparative validation of surgical phase recognition, instrument keypoint estimation, and instrument instance segmentation in endoscopy: Results of the PhaKIR 2024 challenge

### v1

@@ Abstract
Surgical phase recognition, Instrument keypoint estimation, Instrument instance segmentation, Robot-assisted surgery[1]LABEL:#1
1 Introduction
@@
However, the datasets provided by these sub-challenges do not cover all the necessary aspects required for the development of real-world applications. For the phase recognition task, only one paper presents a dataset that originates from multiple centers [13]. In the context of the instrument segmentation task, several challenges do not differentiate between distinct instrument classes [14, 16, 17, 13, 19] or between individual instances within the same class [14, 15, 16, 13, 18, 19]. This lack of distinction hampers the ability to accurately identify overlapping instruments and complicates the task of tracking the tools over time, which relies on the consistent identification of each instrument across successive frames. The recordings of some sub-challenges operate on porcine tissue [15, 16], which significantly simplifies the recognition of surgical instruments compared to human tissue [16], or do not use any biological tissue at all [19]. Some datasets only provide single frames [15, 16, 13, 20] or short clips [17] instead of full video sequences, precluding the use of temporal information for the segmentation task. Information on relevant instrument keypoints is available in only one sub-challenge [14]. However, for in-vivo recordings involving manual instruments, only the coordinates of a single keypoint and the instrument orientation are provided. Meanwhile, sequences featuring robotic instruments are based on ex-vivo data from animal tissue, representing highly simplified movements that do not account for real-world surgical complexities. Additionally, none of the sub-challenge datasets for the segmentation of surgical instruments or the keypoint estimation task include multi-center data, thereby limiting the ability to evaluate the participant methods across different clinical environments and surgical tools.
@@
The appendix contains information regarding the challenge organizers (see A), the data description and labeling instructions (see B), the submission instructions for the participants (see C), as well as the challenge design document (see D).
@@ 5.3.1 Task 1: Surgical phase recognition
For the task of surgical phase recognition, the F1-score [29, 30] was used as the per-class performance metric, while Balanced Accuracy (BA, [31]) served as the metric for evaluating the overall multi-class performance. The F1-score, defined as
represents the harmonic mean of precision and recall, making it particularly suitable for the surgical phase recognition task as it accounts for both false positives and false negatives. Similarly, the BA defined as
@@ 5.3.2 Task 2: Instrument instance segmentation
Here, denotes the precision at a given recall level . The final mAP score is obtained by averaging across all classes and IoU thresholds. Since participants’ algorithms produced only binary segmentation masks without class probabilities, instance matching was based on the IoU between predicted instances and ground truth instances , as defined by
which was originally proposed by [38]. The 95% Hausdorff Distance (HD) quantifies the boundary agreement between two point sets and , typically representing the contours of predicted and ground truth segmentations. It is defined as
@@ 6.1 Task 1: Surgical phase recognition
| augi | LoViT [49] | Yes, unidirectional ASMamba [50] | LoViT pre-trained on Cholec80 [51] | Weighted cross-entropy, phase transition penalties |
@@
| hanglok | TranSVnet [52] | Yes, TeCNO [53] | ResNet-50 backbone pre-trained on Cholec80 [51] | Cross-entropy |
@@
50 (stage 1), 150 (stage 2) /
100 |
| jmees_inc. | EVA-02 [54] | Yes, EMA + post-processing rules | Pre-trained on ImageNet | Cross-entropy |
@@
| ryze | ResNet-50 [55] | Yes, LSTM layer | Pre-trained on ImageNet | Cross-entropy |
@@
| smartlab_hkust | Surgformer [56] | Yes, HTA [56] | Weights of TimeSFormer [57] pre-trained on Kinetics-400 dataset [58] | Cross-entropy |
@@
| uniandes24 | MuST [59] | Yes, MTFE + TCM [59] | MViT [60] backbone pre-trained on Kinetics-400 [58] and Cholec80 [23] | Cross-entropy |
@@
| yipingli | CAFormer [61] | Yes, LSTM layer | Yes, pre-trained weights from SurgeNet [62] | Cross-entropy |
@@ 6.2 Task 2: Instrument instance segmentation
| augi | Mask2Former [70], PINS module | No | Swin-L [71] backbone pre-trained on ImageNet | Binary cross-entropy + dice loss (segmentation), focal loss (classification) |
@@
| goncalo | Mask R-CNN [72] | No | Microsoft COCO [33] | Binary cross-entropy (mask head, objectness), cross-entropy (classification) |
@@
| hanglok | SAM [73], DETR [74] | No | SAM pre-trained ViT-H weights [73], DETR R50 weights [74], DETR backbone pre-trained on ImageNet | Dice loss + NTXentLoss [75, 76] (SAM), cross-entropy + IoU loss (DETR) |
@@
| jmees_inc. | Mask2Former [70], SAM 2 [77] | No | No external data, pseudo labels generated from unlabeled PhaKIR video frames | Cross-entropy (classification), dice loss + cross-entropy (segmentation) |
@@
| kist_harilab | TP-SIS [78], BERT [79], Swin transformer [71] | No | Pre-trained TP-SIS based on CLIP, EV-17, EV-18 [80, 15, 16], BERT pre-trained on BooksCorpus [81] and English wikipedia, Swin backbone pre-trained on ImageNet | Cross-entropy |
@@
| floor9 | RTMDet-Ins-l [69], SAM 2 [77] | Yes, match instances along frames | RTMDet-Ins-l pre-trained on CholecInstanceSeg [82] | Dice-loss |
@@
| uniandes24 | MATIS [85] | Yes, TAPIS [86], SAM2 [77] | Pre-trained ViT on ImageNet, Mask2Former on Microsoft COCO [33], MViT on Kinetics-400 [58], SAM2 on SA-V [77] | Cross-entropy |
@@
Step decay (Mask2Former), cosine annealing (MViT, Transformer) |
@@
In the first stage, feature extraction is performed at multiple scales using a Swin-Large (Swin-L) backbone [71], allowing the model to accurately segment both small and large surgical tools within complex surgical scenes.
@@
| alvaro | YOLOv8 (yolo-v8-x variant) [93] | No | No | Cross-entropy (classification), complete IoU loss (bounding box regression), distributed focal loss (bounding box identification), OKS loss (keypoint localization) |
@@
One cycling lr with cosine annealing |
@@
| sds-hd | YOLOv8 (yolo-v8-x-pose-p6 variant) [93] | No | No | Cross-entropy (classification, keypoint locations and visibility), complete IoU loss (bounding box regression), distributed focal loss (bounding box identification) |
@@ 6.3 Task 3: Instrument keypoint estimation
The following sections outline the methodological approaches employed by the participants in the instrument keypoint estimation task. Table 4 follows the same structure as the Tables 2 and 3 providing a concise overview of each team’s approach. It summarizes the key characteristics of the submitted methods, including the architectures used, the incorporation of temporal information, the utilization of external data, as well as essential training parameters.
@@ 7 Results
The following sections present the evaluation results, beginning with an overview of the two statistical analysis methods used (Sec. 7.1). This is followed by the detailed results for the three challenge tasks: surgical phase recognition (Sec. 7.2), instrument instance segmentation (Sec. 7.3), and instrument keypoint estimation (Sec. 7.4). For each task, we first report the overall performance results (Tab. 5, 7, 11), followed by a statistical analysis of ranking stability (Fig. 5, 9, 14) and and task-specific supplementary analyses. For the phase recognition task, we include a team-specific confusion matrix visualization (Fig. 6). For the segmentation and keypoint estimation tasks, we present qualitative results (Tab. 7.3, 7.3, 7.4). In addition, the instrument segmentation task includes a comparison of the methodological approaches used by the participating teams (Fig. 10). Finally, for all three tasks, we analyze team performance at the level of individual test videos, enabling a more fine-grained assessment of robustness and consistency (Fig. 7, 8, 11, 12, 13, 15; Tab. 6, 10, 13).
7.1 Statistical analysis
@@ 7.2 Task 1: Surgical phase recognition
Table 6. presents quantitative results summarizing the mean metric values for each individual video in the test dataset. These numerical results align well with the KDE visualizations shown in Figures 7 and 8, reinforcing the observed performance trends. Among all test sequences, video nine was he most successfully recognized across teams, achieving an average F1-score of 63.13% and BA of 81.22%. In contrast, video 14 proved most challenging, with the lowest average scores (36.47% F1-score, 68.05% BA).
| Team | DSC (in %) | (in %) | 95% HD | Avg. | Global | ||||||
| 95% CI | Rank | 95% CI | Rank | 95% CI | Rank | rank | rank | ||||
| jmees_inc. | 35.48 | 1 | 36.05 | 1 | 75.48 | 3 | 1.67 | 1st | |||
| uniandes24 | 33.10 | 2 | 33.39 | 3 | 71.47 | 1 | 2.00 | 2nd | |||
| augi | 32.95 | 3 | 33.53 | 2 | 73.10 | 2 | 2.33 | 3rd | |||
| kist_harilab | 26.72 | 4 | 26.40 | 4 | 163.91 | 6 | 4.67 | 4th | |||
| floor9 | 21.64 | 6 | 20.63 | 6 | 82.91 | 4 | 5.33 | 5th | |||
| sk | 24.24 | 5 | 23.73 | 5 | 243.56 | 7 | 5.67 | 6th | |||
| goncalo | 18.57 | 8 | 18.34 | 8 | 155.12 | 5 | 7.00 | 7th | |||
| hanglok | 18.92 | 7 | 18.84 | 7 | 246.13 | 8 | 7.33 | 8th | |||
| recogna | 13.46 | 9 | 12.39 | 9 | 300.12 | 9 | 9.00 | 9th | |||
The second-highest performance was observed for video six (56.80% F1-score, 76.77% BA), followed by video 12 (45.55% F1-score, 72.87% BA), and video eight (44.15% F1-score, 72.30% BA).
@@ 7.4 Task 3: Instrument keypoint estimation
8 Discussion
8.1 Challenge organization
As illustrated in Figure 2, the PhaKIR challenge attracted the highest number of registrations and submissions among all EndoVis-2024 sub-challenges. This underscores both the high relevance of the research questions addressed and the strong interest from the scientific community. It should be noted, however, that PhaKIR comprises three distinct tasks, whereas most other sub-challenges focused on one single task. This difference must be taken into account when interpreting the participation metrics.
@@ 8.7 Overall findings
Finally, cross-hospital comparison of test video results reveals a performance bias: algorithms performed best on videos from hospital 1 (), which was also more prominently represented in the training data. In contrast, performance on videos from hospital 2 () and hospital 3 () was consistently lower. This outcome, while not surprising, highlights a key limitation of all submissions – the challenge of generalizability across unseen clinical environments remains unsolved.

### v2

@@ Abstract
Surgical phase recognition, Instrument keypoint estimation, Instrument instance segmentation, Robot-assisted surgery1 Introduction
@@
However, the datasets provided by these sub-challenges do not cover all the necessary aspects required for the development of real-world applications. For the phase recognition task, only two out of three papers present a dataset that originates from multiple centers [14, 93]. In the context of the instrument segmentation task, several challenges do not differentiate between distinct instrument classes [12, 5, 77, 14, 63] or between individual instances within the same class [12, 6, 5, 14, 70, 63]. This lack of distinction hampers the ability to accurately identify overlapping instruments and complicates the task of tracking the tools over time, which relies on the consistent identification of each instrument across successive frames. The recordings of some sub-challenges operate on porcine tissue [6, 5], which significantly simplifies the recognition of surgical instruments compared to human tissue [5], or do not use any biological tissue at all [63]. Some datasets only provide single frames [6, 5, 14, 13, 63] or short clips [77] instead of full video sequences, precluding the use of temporal information for the segmentation task. Information on relevant instrument keypoints is available in only one sub-challenge [12]. However, for in-vivo recordings involving manual instruments, only the coordinates of a single keypoint and the instrument orientation are provided. Meanwhile, sequences featuring robotic instruments are based on ex-vivo data from animal tissue, representing highly simplified movements that do not account for real-world surgical complexities. Additionally, none of the sub-challenge datasets for the segmentation of surgical instruments or the keypoint estimation task include multi-center data, thereby limiting the ability to evaluate the participant methods across different clinical environments and surgical tools. An overview regarding previous related EndoVis sub-challenges together with the conducted tasks, the number of participating medical centers, and key properties of the challenges and datasets is given in Table 1.
@@
| Challenge | Task | #Med. | Challenge / dataset properties | ||
| PR | IS | KE | centers | ||
| Speidel et al. (2017) [85] | ✓ | 1 | single-centered, single-task, methods no longer considered state of the art | ||
| Wagner et al. (2023) [93] | ✓ | 3 | single-task, strong focus on one leading, more traditional architecture | ||
| Bodenstedt et al. (2021) [14] | ✓ | ✓ | 3 | no instrument classification, no instance segmentation, single frames | |
| Allan et al. (2019) [6] | ✓ | 1 | single-centered, single-task, porcine tissue, no instance segmentation, single frames | ||
| Allan et al. (2020) [5] | ✓ | 1 | single-centered, single-task, porcine tissue, no instrument classification, no instance segmentation, single frames | ||
| Roß et al. (2021) [77] | ✓ | 1 | single-centered, single-task, no instrument classification, short clips | ||
| Malpani and Glock (2023) [63] | ✓ | - | single-centered, single-task, synthetic data, single frames | ||
| Bodenstedt et al. (2023) [13] | ✓ | 1 | single-centered, single-task, single frames | ||
| Psychogyios et al. (2024) [70] | ✓ | 1 | single-centered, single-task, no instance segmentation | ||
| Bodenstedt et al. (2018) [12] | ✓ | ✓ | 1 | single-centered, no instrument classification, no instance segmentation | |
| PhaKIR (2025) | ✓ | ✓ | ✓ | 3 | multi-centered, multi-task, instrument classification, instance segmentation, real-world surgeries, full sequences |
@@ 5.3.1 Task 1: Surgical phase recognition
For the task of surgical phase recognition, the F1-score [91, 19] was used as the per-class performance metric, while Balanced Accuracy (BA, [88]) served as the metric for evaluating the overall multi-class performance. The F1-score, defined as:
represents the harmonic mean of precision and recall, making it particularly suitable for the surgical phase recognition task as it accounts for both false positives and false negatives. Similarly, the BA, defined as:
@@ 5.3.2 Task 2: Instrument instance segmentation
Here, denotes the precision at a given recall level . The final mAP score is obtained by averaging across all classes and IoU thresholds. Since participants’ algorithms produced only binary segmentation masks without class probabilities, instance matching was based on the IoU between predicted instances and ground truth instances , as defined by:
which was originally proposed by [40]. The 95% Hausdorff Distance (HD) quantifies the boundary agreement between two point sets and , typically representing the contours of predicted and ground truth segmentations. It is defined as:
@@ 6.1 Task 1: Surgical phase recognition
| augi | LoViT [56] | Yes, unidirectional ASMamba [16] | LoViT pre-trained on Cholec80 [67] | WCE, phase transition penalties |
@@
| hanglok | TranSVnet [30] | Yes, TeCNO [21] | ResNet-50 backbone pre-trained on Cholec80 [67] | CE |
@@
50 (stage 1),
150 (stage 2) / 100 |
| jmees_inc. | EVA-02 [29] | Yes, EMA + post-processing rules | Pre-trained on ImageNet | CE |
@@
| ryze | ResNet-50 [36] | Yes, LSTM layer | Pre-trained on ImageNet | CE |
@@
| smartlab_hkust | Surgformer [95] | Yes, HTA [95] | Weights of TimeSFormer [10] pre-trained on Kinetics-400 dataset [45] | CE |
@@
| uniandes24 | MuST [68] | Yes, MTFE + TCM [68] | MViT [28] backbone pre-trained on Kinetics-400 [45] and Cholec80 [90] | CE |
@@
| yipingli | CAFormer [97] | Yes, LSTM layer | Yes, pre-trained weights from SurgeNet [41] | CE |
@@ 6.2 Task 2: Instrument instance segmentation
| augi |
Mask2Former [18],
PINS module |
No | Swin-L [57] backbone pre-trained on ImageNet | Binary CE + dice loss (segmentation), focal loss (classification) |
@@
| goncalo | Mask R-CNN [35] | No | Microsoft COCO [55] | Binary CE (mask head, objectness), CE (classification) |
@@
| hanglok |
SAM [48],
DETR [15] |
No | SAM pre-trained ViT-H weights [48], DETR R50 weights [15], DETR backbone pre-trained on ImageNet | Dice loss + NTXentLoss [34, 17] (SAM), CE + IoU loss (DETR) |
@@
| jmees_inc. |
Mask2Former [18],
SAM 2 [72] |
No | No external data, pseudo labels generated from unlabeled PhaKIR video frames | CE (classification), dice loss + CE (segmentation) |
@@
| kist_harilab |
TP-SIS [101],
BERT [23], Swin transformer [57] |
No | Pre-trained TP-SIS based on CLIP, EV-17, EV-18 [71, 6, 5], BERT pre-trained on BooksCorpus [103] and English wikipedia, Swin backbone pre-trained on ImageNet | CE |
@@
| floor9 |
RTMDet-Ins-l [60],
SAM 2 [72] |
Yes, match instances along frames | RTMDet-Ins-l pre-trained on CholecInstanceSeg [4] | Dice-loss |
@@
| uniandes24 | MATIS [7] | Yes, TAPIS [8], SAM2 [72] | Pre-trained ViT on ImageNet, Mask2Former on Microsoft COCO [55], MViT on Kinetics-400 [45], SAM2 on SA-V [72] | CE |
@@
SD (Mask2Former), CA (MViT, Transformer) |
@@
In the first stage, feature extraction is performed at multiple scales using a Swin-Large (Swin-L backbone [57], allowing the model to accurately segment both small and large surgical tools within complex surgical scenes.
@@
| alvaro | YOLOv8 (yolo-v8-x variant) [92] | No | No | CE (classification), complete IoU loss (bounding box regression), distributed focal loss (bounding box identification), OKS loss (keypoint localization) |
@@
OCLR with CA |
@@
| sds-hd | YOLOv8 (yolo-v8-x-pose-p6 variant) [92] | No | No | CE (classification, keypoint locations and visibility), complete IoU loss (bounding box regression), distributed focal loss (bounding box identification) |
@@ 6.3 Task 3: Instrument keypoint estimation
The following sections outline the methodological approaches employed by the participants in the instrument keypoint estimation task. Table 5 follows the same structure as the Tables 3 and 4 providing a concise overview of each team’s approach. It summarizes the key characteristics of the submitted methods, including the architectures used, the incorporation of temporal information, the utilization of external data, as well as essential training parameters.
@@ 7 Results
The following sections present the evaluation results, beginning with an overview of the two statistical analysis methods used (Sec. 7.1). This is followed by the detailed results for the three challenge tasks: surgical phase recognition (Sec. 7.2), instrument instance segmentation (Sec. 7.3), and instrument keypoint estimation (Sec. 7.4). For each task, we first report the overall performance results (Tab. 6, 8, 12), followed by a statistical analysis of ranking stability (Fig. 5, 9, 14) and task-specific supplementary analyses. For the phase recognition task, we include a team-specific confusion matrix visualization (Fig. 6). For the segmentation and keypoint estimation tasks, we present qualitative results (Tab. 7.3, 7.3, 8.1). In addition, the instrument segmentation task includes a comparison of the methodological approaches used by the participating teams (Fig. 10). Finally, for all three tasks, we analyze team performance at the level of individual test videos, enabling a more fine-grained assessment of robustness and consistency (Fig. 7, 8, 11, 12, 13, 15; Tab. 7, 11, 14).
7.1 Statistical analysis
@@ 7.2 Task 1: Surgical phase recognition
| Team | DSC (in %) | (in %) | 95% HD | Avg. | Global | ||||||
| 95% CI | Rank | 95% CI | Rank | 95% CI | Rank | rank | rank | ||||
| jmees_inc. | 35.48 | 1 | 36.05 | 1 | 75.48 | 3 | 1.67 | 1st | |||
| uniandes24 | 33.10 | 2 | 33.39 | 3 | 71.47 | 1 | 2.00 | 2nd | |||
| augi | 32.95 | 3 | 33.53 | 2 | 73.10 | 2 | 2.33 | 3rd | |||
| kist_harilab | 26.72 | 4 | 26.40 | 4 | 163.91 | 6 | 4.67 | 4th | |||
| floor9 | 21.64 | 6 | 20.63 | 6 | 82.91 | 4 | 5.33 | 5th | |||
| sk | 24.24 | 5 | 23.73 | 5 | 243.56 | 7 | 5.67 | 6th | |||
| goncalo | 18.57 | 8 | 18.34 | 8 | 155.12 | 5 | 7.00 | 7th | |||
| hanglok | 18.92 | 7 | 18.84 | 7 | 246.13 | 8 | 7.33 | 8th | |||
| recogna | 13.46 | 9 | 12.39 | 9 | 300.12 | 9 | 9.00 | 9th | |||
@@
Table 7. presents quantitative results summarizing the mean metric values for each individual video in the test dataset. These numerical results align well with the KDE visualizations shown in Figures 7 and 8, reinforcing the observed performance trends. Among all test sequences, video nine was he most successfully recognized across teams, achieving an average F1-score of 63.13% and BA of 81.22%. In contrast, video 14 proved most challenging, with the lowest average scores (36.47% F1-score, 68.05% BA). The second-highest performance was observed for video six (56.80% F1-score, 76.77% BA), followed by video 12 (45.55% F1-score, 72.87% BA), and video eight (44.15% F1-score, 72.30% BA).
@@ 8.1 Challenge organization
| Team | (in %) | ||||
| V_06 | V_08 | V_09 | V_12 | V_14 | |
| sds-hd | 49.79 | 19.58 | 53.08 | 14.17 | 17.46 |
| alvaro | 29.76 | 13.58 | 25.49 | 2.82 | 4.80 |
| mean | 39.78 | 16.58 | 39.28 | 8.49 | 11.13 |
As illustrated in Figure 2, the PhaKIR challenge attracted the highest number of registrations and submissions among all EndoVis-2024 sub-challenges, highlighting both the high relevance of the research questions and the strong interest from the scientific community. Notably, PhaKIR was hosted on a bespoke platform developed according to the principles outlined in the “DIY Challenge Blueprint” publication [49]. Self-hosting provided several advantages that likely contributed to the high level of engagement. The platform was fully customizable to the specific needs of the challenge, enabling streamlined submission workflows, tailored evaluation pipelines, and flexible task definitions. Participants were continuously kept informed via email updates and news posts on the challenge website, reflecting an active communication strategy by the organizers. This close interaction reduced participation barriers and further supported the recruitment of a large and diverse participant cohort. In addition, the large dataset size and the high quality of annotations offered a uniquely rich benchmarking opportunity, enhancing the attractiveness of the challenge. This combination of platform usability, dataset quality, task diversity, and proactive participant engagement distinguished PhaKIR from challenges hosted on standard portals such as Synapse or Grand Challenge. It should be noted, however, that PhaKIR comprises three distinct tasks, whereas most other sub-challenges focused on a single task. This difference should be considered when interpreting participation metrics.
@@ 8.7 Overall findings
Finally, cross-hospital comparison of test video results reveals a performance bias: algorithms performed best on videos from hospital 1 (), which was also more prominently represented in the training data. In contrast, performance on videos from hospital 2 () and hospital 3 () was consistently lower. This results indicate that the predominance of data from hospital 1 within the training set has a significant impact on test performance, enabling the models to achieve improved adaptability to this hospital’s data distribution. This outcome also highlights a key limitation of all submissions – the challenge of generalizability across unseen clinical environments remains unsolved.
@@ 8.8 Future research directions
Additionally, future work could investigate cross-hospital generalization by training models on one or two centers and testing on a completely unseen third. Such experiments would help identify failure modes and provide insights into the robustness of methods under real-world distribution shifts.
Beyond dataset diversity, another key direction is the design of model architectures and algorithms that explicitly promote generalization. Architectural choices, regularization strategies, and training paradigms can all influence a model’s ability to transfer across institutions, equipment, and surgical styles. For instance, approaches that incorporate domain-invariant features, multi-scale representations, or advanced attention mechanisms may improve robustness to distribution shifts. Investigating such strategies in combination with diverse training data could therefore provide a powerful way to enhance real-world applicability of the developed models.